# Multiseed Krylov Complexity

### v1

@@ Abstract
Attractive as it is, the practical performance of Krylov complexity has met some challenges, and our goal here is to present an upgrade that addresses these challenges. A key property one expects from complexity measures of quantum evolution is that they should assign smaller values to integrable/solvable systems than to generic/chaotic systems. If successful, this would give a mathematical expression to the intuitive notion that solved problems are easier than unsolved ones. Krylov complexity tends to grow at early times and saturate at a plateau at late times, and the height of this late-time plateau is one possible indicator of how complex a system is. This program has been seen to work well in some cases, and some general principles have been spelled out for how integrability may reduce the height of the late-time plateau [4]. However, the success of this approach depends on the choice of the initial operator, for which there has been no systematic understanding up to this point. For example, in [5], the performance of the late-time plateau as an indicator of integrability of a spin chain is broken by choosing a particular spatial projection of a single-site spin as the initial seed, while another spatial projection of the same single-site spin leads to the desired performance. Further discussions of the late-time plateau and its dependence on the initial operator can be found in [6, 7, 8, 9].
@@
Lanczos algorithm with single and multiple seeds.— We start with a brief review of the standard single-seed Krylov complexity formulated in terms of the ordinary Lanczos algorithm [19], before proceeding with the multiseed complexity and block Lanczos algorithm. To optimize the notation, we represent every operator $\mathcal{O}$ as a ‘state’ in the space of operators, and write it as $|\mathcal{O}\rangle$. The Krylov basis is constructed given an initial seed operator $|\mathcal{O}_{0}\rangle$ and the Liouvillian $\mathcal{L}=[{H},\cdot]$. It can be defined as an orthogonalization of the Krylov sequence $\big{\{}|\mathcal{O}_{0}\rangle,\mathcal{L}|\mathcal{O}_{0}\rangle,\mathcal{L}^{2}|\mathcal{O}_{0}\rangle,...\big{\}}$ with respect to the inner product $\langle\mathcal{A}|\mathcal{B}\rangle\equiv\operatorname{Tr}[{\mathcal{A}}^{\dagger}{\mathcal{B}}]$. As a consequence of its hermiticity, the Liouvillian is tridiagonal in the Krylov basis $|\mathcal{O}_{j}\rangle$:
@@
In applications, Krylov complexity (3) shows early-time growth, followed by saturation at a plateau at late times. The height of this plateau can be computed [4] as the all-time average of (3), yielding
@@
where $|\omega_{\alpha}\rangle$ are the eigenstates of the Liouvillian $\mathcal{L}$ restricted to the Krylov space. The complexity plateau height has been shown to decrease in the presence of integrability for some systems and for some specific initial operators [4]. However, this link is not general, and it easily breaks down even for relatively simple systems and local operators [5]. We would like to cure this dependence of the Krylov complexity performance on the choice of initial operator.
Inspired by the successes of Nielsen complexity, as mentioned in the introduction, we attempt to stabilize the performance of the Krylov complexity plateau by filling the lowest-weight subspace with all the simplest local operators of the system under consideration (e.g. 1-body operators), and then using the Liouvillian as before to get progressively more complex subspaces. Specifically, we start with a collection of $m$ seed operators, $\big{\{}|\mathcal{O}_{0,0}\rangle,|\mathcal{O}_{0,1}\rangle,...,|\mathcal{O}_{0,m-1}\rangle\big{\}}$, which we can make mutually orthogonal, and which we refer to collectively as ${\Omega}_{0}$. The Krylov basis will now consist of the orthogonalization of all the operators in the set $\big{\{}{\Omega}_{0},{\mathcal{L}}{\Omega}_{0},{\mathcal{L}^{2}}{\Omega}_{0},...\big{\}}$, obtained by repeatedly applying the Liouvillian to the initial set. As before, in finite-dimensional spaces, applying ${\mathcal{L}}$ will eventually yield operators that are all linear combinations of the previous ones, which means the construction of the block Krylov basis $\big{\{}{\Omega}_{0},{\Omega}_{1},...,{\Omega}_{M-1}\big{\}}$ is complete. The operators in ${\Omega}_{J}$ are obtained by orthogonalizing those in ${\mathcal{L}^{J}}{\Omega}_{0}$ against all previous ones and against each other, and normalizing. An operator can be discarded if it is a linear combination of the previous ones, so that ${\Omega}_{J}$ contains $p_{J}$ operators, with $p_{J}$ non-increasing with $J$:
@@
Following the intuition that more applications of the Liouvillian result in more complex operators, we assign complexity $J$ in the range $10,\ldots,M-$ 1 to all operators in ${\Omega}_{J}$, resulting in the following definition of complexity for the evolution of the individual seed operators $\mathcal{O}_{0,n}$:
@@
With this weight assignment, the complexity only depends on the subspace $\Omega_{0}$ spanned by the initial seeds, not on the specific basis chosen within $\Omega_{0}$. With a seed containing $m$ simple local operators in $\Omega_{0}$, as described above, we define multiseed Krylov complexity as the average of (7) over all these seed operators:
@@
Our main interest is in the late-time plateau of this quantity, captured by the all-time average
@@
Numerical implementation.— For any Lanczos-type algorithm — and these issues are more pronounced for block Lanczos algorithms — one needs to manage the numerical instabilities, which otherwise quickly build up at each step due to large subtractions in the course of orthogonalization. In general, this is done by explicitly reorthogonalizing the operators after some number of iterative steps. One can choose to reorthogonalize at every step (full reorthogonalization), or find a way to determine when enough error has accumulated to make reorthogonalization necessary. For our purposes, we have found that even for modest operator space dimensions of a few thousand, guaranteeing orthogonality up to the standard machine precision $\sim\!10^{-16}$ using full reorthogonalization is not enough to get an accurate Krylov basis (the output basis is visibly unstable with respect to increasing the arithmetic precision). We therefore use high-precision arithmetic in order to guarantee orthogonality with as much precision as needed. Furthermore, given the high numerical cost of full reorthogonalization, we opted for the partial reorthogonalization of [21]: after every Lanczos step, one can estimate the current level of orthogonality of the basis, and only reorthogonalize if some threshold has been reached [22]. Specifically, the lists of operators generated in the current and previous step are both reorthogonalized against all previous ones and among themselves. Finally, even for the steps when reorthogonalization is not needed, the current list is orthogonalized against the previous two. If the precision of operations is $\epsilon$, the threshold for reorthogonalization is set to $\sqrt{\epsilon}$, so orthogonality of the final basis is guaranteed only up to $\sqrt{\epsilon}$. Though this approach uses more memory, it is more time-efficient than using full reorthogonalization at precision $\sqrt{\epsilon}$ directly, since it only reorthogonalizes when needed. The rest of the computations involved in calculating (10) do not suffer from instabilities, and were therefore performed at standard precision. We have made all of our numerical scripts public [23].
@@
$N=\sum_{n=0}^{\infty}a_{n}^{\dagger}a_{n},\quad M=\sum_{n=1}^{\infty}n\,a_{n}^{\dagger}a_{n}.\vspace{-1mm}$ (14)
@@
It is natural to consider $k$-body operators simple if $k$ is small. The first choice we make for the seed operators are all 0-body operators: those that leave the occupation numbers of individual modes $a_{n}$ unchanged, as in the Nielsen complexity considerations of [12]. The second option is to choose a smaller set consisting only of the number operators for individual modes: $O_{k}=a_{k}^{\dagger}a_{k}$. The performance of the multiseed Krylov complexity plateau (10) for both seed choices is shown in Fig. 2, where lower values are consistently seen for the integrable case (15).
To summarize, we have developed an upgrade of Krylov complexity as defined in [1] that takes as its initial seed not a single operator but a collection of all simple operators in the theory, selected according to a straightforward physical criterion (for example, all few-body operators). The late-time plateau of this new quantity, given explicitly by (10), reliably assigns lower values to integrable than to chaotic systems within a set of standard test examples typical of quantum chaos considerations. Our construction can be adapted to deal with state complexity, introduced in [3], rather than operator complexity, producing similar results, which we briefly review in the Appendix. There is some apparent similarity between our construction and the notion of ‘operator size’ [33], where one also starts with a set of simplest operators and grades all other operators in order of increasing complexity. Our precise definition differs, however, and takes as its essential input the actual dynamical evolution via the block Lanczos algorithm in order to develop the operator grading, rather than relying on purely lexicographic criteria applied to the operators written out through the elementary degrees of freedom. We show in the Appendix that operator size does not reproduce the successful performance of multiseed Krylov complexity demonstrated above in a series of examples.
@@
Appendix
@@
The operator size is connected to Krylov complexity via the framework of q-complexities introduced in [1]. Other than these two examples, this class of quantities also includes other measures of operator complexity under current investigation, including out-of-time-order correlators (OTOCs). Within this paradigm, Krylov complexity is especially important, as it was shown that it provides an upper bound on the growth of any q-complexity.
The difference between this construction and our multiseed complexity is that the operator size receives no input from the Hamiltonian, and in this sense is purely kinematic rather than dynamical. The entire basis from simple to complex operators has to be specified manually, based for example on the locality or rank of the operators. Since our goal is to characterize the general properties of a given Hamiltonian, we collect the simplest operators of the system and consider their late-time sizes, similarly to what was done in the main text for multiseed Krylov complexity. We start by considering the $m$ simplest operators with $J=1$, and compute each of their sizes under time evolution:
@@
We have observed that this construction, which is the operator size analog of what we did for the multiseed Krylov complexity in the main text, works correctly for Ising chains by assigning smaller complexity values to the integrable case (though the difference between integrable and chaotic cases is very small). By contrast, for the XYZ chain and quantum resonant systems, the operator size fails and assigns higher complexity to integrable cases, as seen in Fig. 4. This shows that, despite some similarities, the successful performance of multiseed Krylov complexity as an integrability measure is not shared by the operator size.

### v2

@@ Abstract
Attractive as it is, the practical performance of Krylov complexity has met some challenges, and our goal here is to present an upgrade that addresses these challenges. A key property one expects from complexity measures of quantum evolution is that they should assign smaller values to integrable/solvable systems than to generic/chaotic systems. If successful, this would give a mathematical expression to the intuitive notion that solved problems are easier than unsolved ones. Krylov complexity tends to grow at early times and saturate at a plateau at late times [4, 5], and the height of this late-time plateau is one possible indicator of how complex a system is. This program has been seen to work well in some cases, and some general principles have been spelled out for how integrability may reduce the height of the late-time plateau [6]. However, the success of this approach depends on the choice of the initial operator, for which there has been no systematic understanding up to this point. For example, in [7], the performance of the late-time plateau as an indicator of integrability of a spin chain is broken by choosing a particular spatial projection of a single-site spin as the initial seed, while another spatial projection of the same single-site spin leads to the desired performance. Further discussions of the late-time plateau and its dependence on the initial operator can be found in [8, 9, 10, 11].
@@
Lanczos algorithm with single and multiple seeds.— We start with a brief review of the standard single-seed Krylov complexity formulated in terms of the ordinary Lanczos algorithm [21], before proceeding with the multiseed complexity and block Lanczos algorithm. To optimize the notation, we represent every operator $\mathcal{O}$ as a ‘state’ in the space of operators, and write it as $|\mathcal{O}\rangle$. The Krylov basis is constructed given an initial seed operator $|\mathcal{O}_{0}\rangle$ and the Liouvillian $\mathcal{L}=[{H},\cdot]$. It can be defined as an orthogonalization of the Krylov sequence $\big{\{}|\mathcal{O}_{0}\rangle,\mathcal{L}|\mathcal{O}_{0}\rangle,\mathcal{L}^{2}|\mathcal{O}_{0}\rangle,...\big{\}}$ with respect to the inner product [22] $\langle\mathcal{A}|\mathcal{B}\rangle\equiv\operatorname{Tr}[{\mathcal{A}}^{\dagger}{\mathcal{B}}]$. As a consequence of its hermiticity, the Liouvillian is tridiagonal in the Krylov basis $|\mathcal{O}_{j}\rangle$:
@@
In applications, Krylov complexity (3) shows early-time growth, followed by saturation at a plateau at late times [4, 5]. The height of this plateau can be computed [6] as the all-time average of (3), yielding
@@
where $|\omega_{\alpha}\rangle$ are the eigenstates of the Liouvillian $\mathcal{L}$ restricted to the Krylov space. (The usage of the all-time average lets one bypass the moment-by-moment dynamical evolution and obtain the late-time saturation value directly.) The complexity plateau height has been shown to decrease in the presence of integrability for some systems and for some specific initial operators [6]. However, this link is not general, and it easily breaks down even for relatively simple systems and local operators [7]. We would like to cure this dependence of the Krylov complexity performance on the choice of initial operator.
Inspired by the successes of Nielsen complexity, as mentioned in the introduction, we attempt to stabilize the performance of the Krylov complexity plateau by filling the lowest-weight subspace with all the simplest local operators of the system under consideration (e.g. 1-body operators), and then using the Liouvillian as before to get progressively more complex subspaces. Specifically, we start with a collection of $m$ seed operators, $\big{\{}|\mathcal{O}_{0,0}\rangle,|\mathcal{O}_{0,1}\rangle,...,|\mathcal{O}_{0,m-1}\rangle\big{\}}$, which we can normalize and make mutually orthogonal, and which we refer to collectively as ${\Omega}_{0}$. The Krylov basis will now consist of the orthogonalization of all the operators in the set $\big{\{}{\Omega}_{0},{\mathcal{L}}{\Omega}_{0},{\mathcal{L}^{2}}{\Omega}_{0},...\big{\}}$, obtained by repeatedly applying the Liouvillian to the initial set. As before, in finite-dimensional spaces, applying ${\mathcal{L}}$ will eventually yield operators that are all linear combinations of the previous ones, which means the construction of the block Krylov basis $\big{\{}{\Omega}_{0},{\Omega}_{1},...,{\Omega}_{M-1}\big{\}}$ is complete. The operators in ${\Omega}_{J}$ are obtained by orthogonalizing those in ${\mathcal{L}^{J}}{\Omega}_{0}$ against all previous ones and against each other, and normalizing. An operator can be discarded if it is a linear combination of the previous ones, so that ${\Omega}_{J}$ contains $p_{J}$ operators, with $p_{J}$ non-increasing with $J$:
@@
Following the intuition that more applications of the Liouvillian result in more complex operators, we assign complexity $J$ in the range $10,\ldots,M-$ 1 to all operators in ${\Omega}_{J}$, resulting in the following definition [24] of complexity for the evolution of each individual seed operator $\mathcal{O}_{0,n}$ in $\Omega_{0}$ indexed by $n$:
@@
As for the standard Krylov complexity, one can think of (7) as an ‘average position’ of the evolving operator, but now relative to the block-Lanczos iterative level $J$, rather than to the ordinary Krylov basis numbering. With a seed containing $m$ simple local operators in $\Omega_{0}$, as described above, we define multiseed Krylov complexity as the average of (7) over all these seed operators:
@@
With the weight assignment in (7), the complexity only depends on the subspace $\Omega_{0}$ spanned by the initial seeds, not on the specific basis chosen within $\Omega_{0}$. Our main interest is in the late-time plateau of this quantity, captured by the all-time average
@@
Numerical implementation.— For any Lanczos-type algorithm — and these issues are more pronounced for block Lanczos algorithms — one needs to manage the numerical instabilities, which otherwise quickly build up at each step due to large subtractions in the course of orthogonalization. In general, this is done by explicitly reorthogonalizing the operators after some number of iterative steps. One can choose to reorthogonalize at every step (full reorthogonalization), or find a way to determine when enough error has accumulated to make reorthogonalization necessary. For our purposes, we have found that even for modest operator space dimensions of a few thousand, guaranteeing orthogonality up to the standard machine precision $\sim\!10^{-16}$ using full reorthogonalization is not enough to get an accurate Krylov basis (the output basis is visibly unstable with respect to increasing the arithmetic precision). We therefore use high-precision arithmetic in order to guarantee orthogonality with as much precision as needed. Furthermore, given the high numerical cost of full reorthogonalization, we opted for the partial reorthogonalization of [25]: after every Lanczos step, one can estimate the current level of orthogonality of the basis, and only reorthogonalize if some threshold has been reached [26]. Specifically, the operators generated in the current and previous steps are both reorthogonalized against all previous ones and among themselves. Finally, even for the steps when reorthogonalization is not needed, the current list is orthogonalized against the previous two. If the precision of operations is $\epsilon$, the threshold for reorthogonalization is set to $\sqrt{\epsilon}$, so orthogonality of the final basis is guaranteed only up to $\sqrt{\epsilon}$. Though this approach uses more memory, it is more time-efficient than using full reorthogonalization at precision $\sqrt{\epsilon}$ directly, since it only reorthogonalizes when needed. The rest of the computations involved in calculating (10) do not suffer from instabilities, and were therefore performed at standard precision. We have made all of our numerical scripts public [27].
@@
$N=\sum_{n=0}^{\infty}a_{n}^{\dagger}a_{n},\quad M=\sum_{n=1}^{\infty}n\,a_{n}^{\dagger}a_{n}.$ (14)
@@
It is natural to consider $k$-body operators simple if $k$ is small. The first choice we make for the seed operators are all 0-body operators: those that leave the occupation numbers of individual modes $a_{n}$ unchanged, as in the Nielsen complexity considerations of [14]. The second option [37] is to choose a smaller set consisting only of the number operators for individual modes: $O_{k}=a_{k}^{\dagger}a_{k}$. The performance of the multiseed Krylov complexity plateau (10) for both seed choices is shown in Fig. 2, where lower values are consistently seen for the integrable case (15).
To summarize, we have developed an upgrade of Krylov complexity as defined in [1] that takes as its initial seed not a single operator but a collection of all simple operators in the theory [38], selected according to a straightforward physical criterion (for example, all few-body operators). The late-time plateau of this new quantity, given explicitly by (10), reliably assigns lower values to integrable than to chaotic systems within a set of standard test examples typical of quantum chaos considerations. Our construction can be adapted to deal with state complexity, introduced in [3], rather than operator complexity, producing similar results, which we briefly review in the supplemental material. There is some apparent similarity between our construction and the notion of ‘operator size’ [39], where one also starts with a set of simplest operators and grades all other operators in order of increasing complexity. Our precise definition differs, however, and takes as its essential input the actual dynamical evolution via the block Lanczos algorithm in order to develop the operator grading, rather than relying on purely lexicographic criteria applied to the operators written out through the elementary degrees of freedom. We show in the supplemental material that operator size does not reproduce the successful performance of multiseed Krylov complexity demonstrated above in a series of examples.
@@ Acknowledgements.
SUPPLEMENTAL MATERIAL
Preferred role of the block Krylov basis
In [3], it was shown that the Krylov basis minimizes a cost function quantifying the spread of the initial state over a given basis. We can use similar arguments in the operator (rather than state) complexity context to show that the block Krylov basis minimizes a different but related cost function. In the following, we give the necessary definitions before stating the theorem and giving the proof.
The block Krylov basis $\mathcal{K}_{B}$ corresponding to an initial seed $\Omega_{0}=\big{\{}|\mathcal{O}_{0,0}\rangle,|\mathcal{O}_{0,1}\rangle,...,|\mathcal{O}_{0,m-1}\rangle\big{\}}$ consists of a sequence of subspaces $\Omega_{J}$, each having dimension $p_{J}$. Using these $p_{J}$, we can take any complete ordered orthogonal basis of operators $\mathcal{B}$ and define a cost function as follows: assign the first $p_{0}$ vectors a weight of $c_{0}$, the next $p_{1}$ a weight of $c_{1}$ and so on, with $c_{J}$ positive and increasing. Since this prescription will not necessarily cover a complete basis, we assign the final weight $c_{M-1}$ to every remaining operator in $\mathcal{B}$ instead of just the next $p_{M-1}$. We denote the subspace with weight $c_{J}$ as $B_{J}$, and the individual elements of $B_{J}$ as $|B_{J,k}\rangle$. The cost function is then defined by:
$\mathcal{C}_{\mathcal{B}}(t)\!=\!\frac{1}{m}\sum_{J,k,n}c_{J}|\langle\mathcal{O}_{0,n}(t)|B_{J,k}\rangle|^{2}\!=\!\sum_{J}c_{J}p_{\mathcal{B}}(J,t),$ (17)
with
$p_{\mathcal{B}}(J,t)=\frac{1}{m}\sum_{k,n}|\langle\mathcal{O}_{0,n}(t)|B_{J,k}\rangle|^{2}$ (18)
representing the average probability, over all elements of $\Omega_{0}$, to be in subspace $J$ at time $t$. Intuitively, this cost function quantifies the spread of the initial subspace of operators $\Omega_{0}$ into the sequence of subspaces $B_{J}$ at time $t$.
Following [3], we consider the sequence of derivatives of $\mathcal{C}_{\mathcal{B}_{1}}(t)$ and $\mathcal{C}_{\mathcal{B}_{2}}(t)$ evaluated at $t=0$, for two different bases $\mathcal{B}_{1}$, $\mathcal{B}_{2}$. If the first inequality in this sequence is $\mathcal{C}^{(l)}_{\mathcal{B}_{1}}(0)<\mathcal{C}^{(l)}_{\mathcal{B}_{2}}(0)$ then $\mathcal{C}_{\mathcal{B}_{1}}(t)<\mathcal{C}_{\mathcal{B}_{2}}(t)$ for any $t$ close enough to 00. So a functional minimization, valid only at early times, can be defined by comparing the sequence of derivatives at $t=0$.
Equation (17) is invariant with respect to changes between orthonormal bases if the subspaces $B_{J}$ remain the same. Given this, consider some complete basis $\mathcal{B}$ for which the $B_{J}$ are the same as the block Krylov subspaces $\Omega_{J}$. $\mathcal{B}$ is complete, so the final subspace $B_{M-1}$ will generically contain the corresponding block Krylov subspace $\Omega_{M-1}$, along with some extra elements; the latter are unexplored by the time evolution of $\Omega_{0}$ since this is completely covered by the block Krylov basis, so they do not affect the final cost. Therefore, we have that $\mathcal{C}_{\mathcal{B}}(t)=\mathcal{C}_{\mathcal{K}_{B}}(t)$, allowing us to identify any such basis as equivalent to the block Krylov basis $\mathcal{K}_{B}$.
Theorem: The block Krylov basis $\mathcal{K}_{B}$, or any equivalent basis, minimizes the cost function (17), with ‘minimization’ understood in the sense described above.
Proof: The proof follows that of [3], but we examine the first $N$ subspaces instead of $N$ elements of the basis. The subspace occupied at $t=0$ is by definition $\Omega_{0}$. A basis whose first subspace is equal to this, $B_{0}=\Omega_{0}$, will have cost $\mathcal{C}^{(0)}_{\mathcal{B}}(0)=c_{0}$. For any other basis, part of the initial subspace $\Omega_{0}$ does not appear in $B_{0}$, so it must appear in some of the later subspaces. This will cause the sum of (17) to receive contributions from the higher weights, which means the cost for such a basis will be higher at $t=0$.
We can therefore assume the basis $\mathcal{B}$ is such that the first $N>0$ subspaces coincide with those of the block Krylov basis, $B_{J}=\Omega_{J}$ for $J=0,1,...,N-1$. If $N=M$, all subspaces coincide and $\mathcal{B}$ has the same cost as $\mathcal{K}_{B}$. Similarly, if $N=M-1$, the next subspace $B_{M-1}$ will be assigned cost $c_{M-1}$ and will contain the rest of the Hilbert space, including the last remaining block Krylov subspace $\Omega_{M-1}$, so again this will have the same cost as $\mathcal{K}_{B}$. So we can assume $N<M-1$, and that $B_{N}\neq\Omega_{N}$. We can expand the time derivatives of the probabilities appearing in (17) using $\partial_{t}|O(t)\rangle=i\mathcal{L}|O(t)\rangle$ as follows:
$p^{(l)}_{\mathcal{B}}(J,t)=\partial_{t}^{l}p_{\mathcal{B}}(J,t)=\frac{i^{l}}{m}\sum_{k,n}\sum_{j=0}^{l}(-1)^{l-j}\binom{l}{j}$
×$\hskip 28.45274pt\times\langle\mathcal{O}_{0,n}(t)|\mathcal{L}^{l-j}|B_{J,k}\rangle\langle B_{J,k}|\mathcal{L}^{j}|\mathcal{O}_{0,n}(t)\rangle.$ (19)
For $J<N$, only the subspace $B_{J}=\Omega_{J}$ appears in the expression for $p^{(l)}_{\mathcal{B}}(J,t)$. It follows that $p^{(l)}_{\mathcal{B}}(J,t)=p^{(l)}_{\mathcal{K}_{B}}(J,t)$ for all $l$ for $J<N$. Next, we focus on the first $12N-$ 1 derivatives and look at the contributions from different $J$.
Evaluating (19) at $t=0$, we get terms containing $\mathcal{L}^{j}|\mathcal{O}_{0,n}(0)\rangle$. For $j<N$, such a term must be entirely contained in the subspaces $B_{J}=\Omega_{J}$ with $J<N$, since this is how the block Krylov subspaces are defined. So we must have $\langle B_{J,k}|\mathcal{L}^{j}|\mathcal{O}_{0,n}(0)\rangle=0$ for all $J\geq N$ and $j<N$. Because we are considering the first $12N-$ 1 derivatives, we also have that $l<2N$, so for any integer $j$ we have that either $j<N$ or $l-j<N$. Each term in the sum (19) at $t=0$ is proportional to either $\langle B_{J,k}|\mathcal{L}^{j}|\mathcal{O}_{0,n}(0)\rangle$ or $\langle B_{J,k}|\mathcal{L}^{l-j}|\mathcal{O}_{0,n}(0)\rangle$, so it vanishes if $J\geq N$ and $l<2N$. Combining the results we have so far, $p^{(l)}_{\mathcal{B}}(J,t)=p^{(l)}_{\mathcal{K}_{B}}(J,t)$ for all $J$ and any $l<2N$. To see which basis minimizes the cost function at early times, we therefore have to examine the next derivative.
Taking $l=2N$, we can again use the fact that $B_{J}=\Omega_{J}$ for $J<N$ to eliminate most of the terms in (19) at $t=0$. The only remaining term corresponds to $j=l-j=N$:
$p^{(2N)}_{\mathcal{B}}(J,0)=\frac{1}{m}\sum_{k,n}\binom{2N}{N}$
×$\hskip 28.45274pt\times\langle\mathcal{O}_{0,n}(0)|\mathcal{L}^{N}|B_{J,k}\rangle\langle B_{J,k}|\mathcal{L}^{N}|\mathcal{O}_{0,n}(0)\rangle.$ (20)
There can be a component of $\mathcal{L}^{N}|\mathcal{O}_{0,n}(0)\rangle$ which is orthogonal to all the subspaces $B_{0},...,B_{N-1}$, call it $|X_{n}\rangle$. Because $B_{N}\neq\Omega_{N}$, we must have $\langle X_{n}|X_{n}\rangle>0$ for some $n$. Since the basis $\mathcal{B}$ is orthogonal, the only contributions to the probability that can appear for $J\geq N$ are from the $|X_{n}\rangle$:
$p^{(2N)}_{\mathcal{B}}(J\geq N,0)=\frac{1}{m}\sum_{k,n}\binom{2N}{N}\langle X_{n}|B_{J,k}\rangle\langle B_{J,k}|X_{n}\rangle.$ (21)
Given that for $J<N$ the probabilities are equal to those of the block Krylov basis, we have every term we need in order to compare the $2N$ derivatives of the cost function:
$\mathcal{C}$ ${}_{\mathcal{B}}^{(2N)}(0)=\sum_{J}c_{J}p_{\mathcal{B}}^{(2N)}(J,0)=\sum_{J=0}^{N-1}c_{J}p_{\mathcal{B}}^{(2N)}(J,0)$
$+\frac{1}{m}\sum_{k,n}\sum_{J=N}^{M-1}\binom{2N}{N}c_{J}\langle X_{n}|B_{J,k}\rangle\langle B_{J,k}|X_{n}\rangle.$ (22)
In the first sum, we can replace all derivatives of the probabilities with the corresponding block Krylov ones, since $J<N$. Furthermore, in the second term, we can replace all the $c_{J}$ with just $c_{N}$ to get the following inequality:
$\mathcal{C}_{\mathcal{B}}^{(2N)}(0)$ $\geq\sum_{J=0}^{N-1}c_{J}p_{\mathcal{K}_{B}}^{(2N)}(J,0)$
$+\frac{c_{N}}{m}\binom{2N}{N}\sum_{k,n}\sum_{J=N}^{M-1}\langle X_{n}|B_{J,k}\rangle\langle B_{J,k}|X_{n}\rangle.$ (23)
The only way to get equality is when only $c_{N}$ contributes to the second sum of (22). This happens only if all the $|X_{n}\rangle$ are contained in the subspace $B_{N}$, which is not the case because we are assuming $B_{N}\neq\Omega_{N}$. We therefore have a strict inequality. We have already argued that the $|X_{n}\rangle$ only have support in $B_{J}$ for $J\geq N$, so in the last term of (23), the sums over $J,k$ represent the full inner product of each $|X_{n}\rangle$:
$\mathcal{C}_{\mathcal{B}}^{(2N)}(0)>\sum_{J=0}^{N-1}c_{J}p_{\mathcal{K}_{B}}^{(2N)}(J,0)+\frac{c_{N}}{m}\binom{2N}{N}\sum_{n}\langle X_{n}|X_{n}\rangle.$ (24)
These are also the terms that appear in $\mathcal{C}_{\mathcal{K}_{B}}^{(2N)}(0)$, since all the $|X_{n}\rangle$ are by definition contained in $\Omega_{N}$, and will be assigned weight $c_{N}$. So we have:
$\mathcal{C}_{\mathcal{B}}^{(2N)}(0)>\mathcal{C}_{\mathcal{K}_{B}}^{(2N)}(0),$ (25)
which means the cost function for $\mathcal{B}$ is greater than that for $\mathcal{K}_{B}$ at early times.
@@
The operator size is connected to Krylov complexity via the framework of q-complexities introduced in [1]. Other than these two examples, this class of quantities also includes other measures of operator complexity under current investigation, including out-of-time-order correlators (OTOCs). Within this paradigm, Krylov complexity is especially important, as it was shown that it provides an upper bound on the growth of any q-complexity. (Note that this, in particular, implies that standard Krylov complexity would provide an upper bound on the multiseed complexity evaluated for individual seed operators. This upper bound does not, however, share the successful performance of the full multiseed complexity for distinguishing integrability and chaos, as evident from our exposition in the main text.)
The difference between this construction and our multiseed complexity is that the operator size receives no input from the detailed specification of the Hamiltonian, and in this sense is purely kinematic rather than dynamical. The entire basis from simple to complex operators has to be specified manually, based for example on the locality or rank of the operators. Since our goal is to characterize the general properties of a given Hamiltonian, we collect the simplest operators of the system and consider their late-time sizes, similarly to what was done in the main text for multiseed Krylov complexity. We start by considering the $m$ simplest operators with $J=1$, and compute each of their sizes under time evolution:
@@
We have observed that this construction, which is the operator size analog of what we did for the multiseed Krylov complexity in the main text, works correctly for Ising chains by assigning smaller complexity values to the integrable case (though the difference between integrable and chaotic cases is very small). By contrast, for the XYZ chain and quantum resonant systems, the operator size fails and assigns higher complexity to the integrable cases, as seen in Fig. 4. This shows that, despite some similarities, the successful performance of multiseed Krylov complexity as an integrability measure is not shared by the operator size.
Averaged single-seed complexity
QRSXYZIsingaveraged $\overline{C_{K}}$
Figure 5: Normalized ordinary Krylov complexity plateau averaged over 1-body operators for Ising and XYZ chains, both with $L=4$, and for quantum resonant systems (QRS) with $N=M=9$. For each integrable-chaotic (white-black) pairing, we divide both results by the highest iterative level reached by the Lanczos algorithm (maximized over all seeds, trials, and jointly maximized over the integrable and chaotic cases), which puts all the heights into the interval $(0,1)$.
In the definition of multiseed complexity, we developed a grading of operators based on the block Lanczos algorithm (qualitatively, in order of increasing complexity), then introduced the quantum-mechanical average of the position with respect to this graded basis for trajectories starting with individual operators at the lowest (simplest) level, and then averaged over all simple initial operators. It is a natural thought that one could apply the same construct to the original Krylov complexity of [1], computing it for each simple operator and then averaging over all such operators.
Before exploring this quantity any further, we remark that it suffers from a number of considerable disadvantages. First, it is much more numerically demanding that our multiseed construction. In the multiseed construction, one runs the (block) Lanczos algorithm only once. For the averaged Krylov complexity, one will have to rerun it for each operator in the simple set, which is a huge number of runs. For instance, the number of 1-body operators is, roughly, the number of degrees of freedom, which by definition diverges in the thermodynamic limit.
Even more dramatically, the Krylov complexity average is sensitive to the choice of basis in the space of simple operators, which is not the case for multiseed complexity. The latter comes from two key properties: first, the subspaces generated by the block Lanczos algorithm depend only on the span of the initial seed, not on the specific basis; second, the multiseed complexity also only depends on these subspaces. Specifically, the expression involves a sum of projections over some basis within each subspace, ensuring that the result is basis-independent. This appealing property of multiseed complexity (which resolves some of the ambiguity inherent in standard Krylov complexity) is not shared by the average of the single-seed Krylov complexity (which would require an explicit basis specification as its input). For example, we found that simply rotating the $(x,y,z)$ axes for the seed operators in a spin chain, while not affecting the multiseed complexity, can drastically alter the averaged Krylov complexity.
Despite these flaws of the averaged-Krylov construction, we find it useful to report the results here, for the relatively small systems we can access numerically: the averaged Krylov complexity plateau distinguishes integrable and chaotic cases correctly for spin chains, but not for quantum resonant systems, as seen in Fig. 5. (As our simple operator basis we took the explicit sets of operators described in our considerations of multiseed complexity, with the same spatial axes used in the expressions for the Hamiltonians.)